### v1

# Cryptogenic Stroke and Migraine: Using Probabilistic Independence and Machine Learning to Uncover Latent Sources of Disease from the Electronic Health Record

Abstract
Migraine is a common but complex neurological disorder that doubles the lifetime risk of cryptogenic stroke (CS). However, this relationship remains poorly characterized, and few clinical guidelines exist to reduce this associated risk. We therefore propose a data-driven approach to extract probabilistically-independent sources from electronic health record (EHR) data and create a 10-year risk-predictive model for CS in migraine patients. These sources represent external latent variables acting on the causal graph constructed from the EHR data and approximate root causes of CS in our population. A random forest model trained on patient expressions of these sources demonstrated good accuracy (ROC 0.771) and identified the top 10 most predictive sources of CS in migraine patients. These sources revealed that pharmacologic interventions were the most important factor in minimizing CS risk in our population and identified a factor related to allergic rhinitis as a potential causative source of CS in migraine patients.
Introduction
Ischemic stroke (IS) is consistently among the top five most common causes of death both in the United States[1] as well as worldwide[2] and is a source of significant health-care related costs in developed countries[3]. Every year, significant resources are spent in the prevention of the ”big three” causes of IS—large vessel atherosclerotic, small-vessel occlusive (lacunar), and cardioembolic [4, 5]. However, much less light has been shed on risk-reduction of cryptogenic stroke (CS), which by some estimates account for 30%-40% of IS cases in the U.S[6]. While the simplest definition of CS includes any ischemic stroke that is not determined to be one of the three aforementioned etiologies[7] (the definition we use in this paper), it should be noted that terminology surrounding CS varies. Some sources use the label of CS to refer only to strokes of undetermined etiology,[8, 9] while others also include strokes of determined but uncommon etiologies[10] and/or embolic strokes of undetermined source (ESUS)[11] under the cryptogenic umbrella.
Despite CS’s relative lack of characterization, a link between migraine and CS has been frequently described in the literature [12]. Migraine is exceedingly common, affecting 10-20% of US residents,[13] and increases the lifetime CS risk by a factor of two. [12] However, few clinical recommendations exist in the literature to reduce this associated stroke risk in migraine patients. These range from from the well-established guideline of avoiding estrogen-containing contraception in patients who experience migraine with aura[14, 15] to identification and closure of a patent foramen ovale [16, 17]—a common procedure that has become controversial in its off-label use in addressing migraine symptoms[18].
Given the poor outcomes associated with stroke and the high prevalence of migraine, identification of additional interventions to reduce the risk of CS in migraine presents a cost-efficient opportunity to reduce overall population morbidity and mortality. The electronic health record (EHR) has long been considered as a potential source of data for investigations into such associations, but has historically been considered too incomplete, noisy, and massive to analyze with typical methods[19, 20]. Methods utilizing machine learning (ML) have already shown success in creating models from analyzing messy, irregular EHR data[19, 21, 20, 22], though these have historically been tuned for predictive accuracy [23, 24] as opposed to causal analysis[25], limiting interpretability. So while current methods of analysis may yield highly accurate predictive models, they tend to not offer insight into the driving factors behind the disease itself[25]. However, recent studies have demonstrated the ability of probabilistic independence-based algorithms[26, 22] to create models that successfully identify patient-specific root causes of disease[27] with only a mild impact on predictive accuracy[22]. These causative models exhibit enhanced interpretability and allow for identification of complex diseases with heterogenous presentations and identification of their driving factors[25, 28].
In this study, we make the following contributions. First, we use probabilistic independence-based techniques to disentangle 2000 latent sources of disease and their clinical signatures from the real-world EHR data of a broad sample of neurology patients. Second, we project the EHR data of a large sample of migraine patients onto this signature space use these projections to train a causal model that predicts a patient’s 10-year risk of CS. Third, we evaluate this model using a held-out test set to assess its accuracy and compare it to alternative models. Finally, we identify which sources have the largest causal effects on CS in migraine patients and may serve as potential targets for future investigation.
Methods
This study was conducted at Vanderbilt University Medical Center (VUMC) and was determined by the Vanderbilt University Institutional Review Board to be non-human-subjects research (#171392).
Table 1: Demographic data of study populations and subsets. IQR: interquartile range
Evaluation Set
(Migraine patients)
Demographic Discovery Set
Negative
for CS
Positive
for CS
N of unique records
309,759 71,206 1,670
Male 46.1% 23.4% 25.5%
Sex Female 53.9% 76.6% 74.4%
White 74.9% 78.0% 73.8%
Black 12.2% 10.6% 16.5%
Other 5.2% 4.9% 2.9%
Race Unknown 9.0% 7.8% 3.1%
Mean 48 42 51
Age, yrs IQR [32, 66] [28, 57] [40, 67]
Mean 9 10 12
Record length, yrs IQR [3, 15] [3, 16] [4, 20]
Data collection and set definitions
All data was extracted from the Synthetic Derivative (SD) [29], VUMC’s de-identified EHR mirror of over 3 million patient charts from 2004-present. With a cutoff date of 08/31/24, we collected historical EHR data across the 9000 most common quantitative clinical variables from all 311,429 patient records that contained at least one ICD-10 code in the range of G40-G47, representing a broad range of paroxysmal neurological disease. We then defined our evaluation set as the subset of these records that contain an ICD-10 code of G43, corresponding to any migraine diagnosis ($N=72,876$). Each patient ID in the evaluation set was then randomly assigned to either the training set ($N=54,654$) or the test set ($N=18,222$) with a probability of 0.2 of being assigned to the latter. Patient IDs in the test set were then subtracted from our initial study population to ensure there was no data leak from patient records in the test set during model training, finalizing our discovery set ($N=309,759$) —demographics detailed in Table 1.
Data pre-processing and continuation
Instances of a patient’s health data occur at discrete, asynchronous points in the record, with large gaps existing between encounters with the healthcare system. Sampling a record at a random point is thus unlikely to give values for a majority of chart variables. This noisy, discrete data must be transformed into continuous and synchronous data that can be sampled at any point in time. We followed the methods employed by Lasko et al.[26] and Mota et al.[22] for data preprocessing. In summary, we computed longitudinal curves for all 9000 chart variables for each patient with a temporal resolution of one day from the discrete and asynchronous EHR observations (Fig.1). The process used for curve generation for each variable depended on its data modality.
For medical conditions, represented by SNOMED billing codes in the SD, we constructed smooth code-intensity curves from discrete instances using a variation of the Random Average Shifted Histograms (RASH) algorithm[30]. In the event of missing data, we imputed a constant curve with baseline annual event-arrival probability of 1/20. For clinical measurements such as vital signs and laboratory values, we used the Piecewise Cubic Hermite Interpolating Polynomial (PCHIP)[31] to generate smooth continuous curves between the recorded discrete values. Records without observations for a given variable were assigned a constant curve of the population median. For medications, we opted to create binary curves that approximate medication adherence, given that medication data in the SD solely indicates presence on a visit’s medication list. We assumed continuation of any regimen if noted on contiguous visits, imputing a value of 1. If a subsequent visit’s list dropped a medication, we assumed discontinuation at the temporal midpoint. Medications that were not noted at all on a given chart were assigned the zero function. Finally, demographics such as race and sex were represented by constant binary variables based on the record’s most recent values.
Figure 1: Preprocessing pipeline for EHR data. Longitudinal curves are inferred from the raw observational data and stacked, synchronizing the discrete data channels and allowing for all clinical variables to be sampled at any given point in time. Random instances of these curvesets records are then sampled and concatenated into the stacked data matrix X.
We stacked each curve longitudinally into an $m\times t$ curveset for each patient, where $m$ is the number of clinical variables and $t$ is the number of days spanned by the record. We then sampled $n$ curveset cross sections at uniformly random times with an average density of one sample per record-year. This results in longer charts having a higher probability of being sampled more than once, while some shorter charts were not sampled at all.
Merging all cross sections yields a dense data matrix $\textbf{X}\in\mathbb{R}^{m\times n}$ where each column $\mathbf{x}_{j}$ of X represents a single estimate of a complete patient state taken at a specific time in a patients record across all $m=9000$ extracted clinical variables. Our matrix X was then standardized by row to bring all variables into roughly the same scale; measurements, log-transformed condition code intensities, and ages were centered so that their means were zero and scaled by two standard deviations[32] while binary variables were left untouched.
Clinical signature discovery
[…]
The FastICA algorithm ensures the rows of S are (approximately) mutually independent; thus each row $\mathbf{s}^{\prime}_{i}$ represents the expression level of a disentangled latent source of disease. Each column $\mathbf{s}_{j}$ of S then quantifies the extent to which each source $j$ is expressed in patient record instance $\mathbf{x}_{j}$. Meanwhile, each column $\mathbf{a}_{j}$ of A represents the clinical signature of source $j$ —i.e. the $m$ linear mixing weights that describe the unique imprint of changes to the EHR data by its corresponding latent source $\mathbf{s}^{\prime}_{j}$. Note that $k$, the number of sources inferred by our ICA decomposition, would ideally be close to the intrinsic dimensionality of X to avoid over- or under-learning [33]; however, we were limited to a value of $k=2000$ due to memory constraints.
Record labeling
Due to cryptogenic stroke’s nature as a diagnosis of exclusion and the varying definitions that get attached to it, there is unsurprisingly no single ICD-10 code or even set of ICD-10 codes that neatly correspond to a diagnosis of CS. Additionally, some non-specific diagnosis codes may be applied to both cryptogenic and non-cryptogenic strokes, and clinicians may sometimes use inaccurate billing codes when documenting a stroke event. To account for these complexities, we developed a 4-step algorithm of inclusion and exclusion criteria in order to label patient records in our evaluation set as positive ($N=1670$) or negative ($N=71,206$) for cryptogenic stroke (Fig. 2).
First, the set of ICD-10 stroke codes were identified that could feasibly be applied to an episode of CS to form an initial set of inclusion criteria (G43.6, I63, I67.8). These codes were then evaluated for how often they coincided with other ICD-10 codes for IS that are specific for a non-cryptogenic etiology in the same patient chart. Codes that had more than a 30% coincidence rate were labelled as overly non-specific for CS and removed as an inclusion criterion, yielding a more specific set of inclusion codes (G43.6, I63.212, I63.52, I63.6, I63.8, I63.9, I67.848). The same stroke codes that are specific for a non-cryptogenic etiology were then used as exclusion criteria to account for patients who initially received a diagnosis of cryptogenic stroke but whose stroke was eventually determined to be of another etiology (I63.0, I63.1, I63.3, I63.4). Additionally, chart review revealed that in the majority of records with only one instance of a stroke code, the code was either applied erroneously or was applied to cases of suspected stroke that were later determined to be negative after further workup. Therefore, having an instance of only one stroke code was chosen as a final exclusion criterion.
[…]
Figure 2: Flowchart detailing the multi-step inclusion and exclusion criteria (based on ICD-10 diagnosis) codes used to label patient records as either positive or negative for CS.
Note that the number of records labeled as positive for CS in our evaluation set is roughly 40% of the number of stroke patients in our evaluation set. This proportion is in line with current literature estimates,[6, 4] suggesting that these criteria are not grossly over- or under-counting instances of cryptogenic stroke.
Model creation
The longitudinal curves for all patient charts in the evaluation set were sampled (Fig.1) with an average density of one sample per record-year within a 10-year sampling window. For patient records without any instance of a stroke diagnosis code, the end of this window coincides with the end of the record. For patient records with a stroke code, the window ends one month prior to the earliest occurrence of a stroke code in order to leave a buffer such that data from the encounter where the stroke was diagnosed is excluded. This results in a prediction time horizon of one month to 121 months; in other words, the model is predicting whether a patient will have a cryptogenic stroke within the next 10 years. Notably, we modified the sampling distribution to ensure that exactly one sample fell within the final year of each record’s window while maintaining our density of one sample per record-year. While this results in shorter records being disproportionately represented in the model, it ensures we have a wide variety of patient records on which to train and test the model, leading to better generalizability.
Just as was done with the discovery set data, each of these synchronized patient chart instances $\mathbf{e}_{j}$ were then stacked longitudinally into a dense data matrix E. This matrix was then projected onto the clinical signature space by multiplying by the inverse of the mixing matrix discovered by our earlier ICA decomposition: $\textbf{S}=\textbf{A}^{-1}\textbf{E}$. Each pair of corresponding columns $\mathbf{e}_{j}$ in E and $\mathbf{s}_{j}$ in S are therefore representations of the same information but in two different vector bases, with the latter encoding each instance of patient data how strongly it expresses each of the 2000 learned signatures. We then used these projections, along with their corresponding labels (Fig. 1(b)) to train a non-linear random forest (RF) model (scikit-learn 1.6.0, Python 3.10.16) to predict 10-year cryptogenic stroke risk. Although alternative architectures such as XGBoost are generally more accurate when making predictive models, [21] RF models have shown superior accuracy and stability in estimating causal effects [26, 22], leading to this study’s choice of model architecture. The optuna package was used to search for the optimal model hyperparameters with the 6-fold cross-validation area under the receiver operating characteristic curve (AUROC) as the primary endpoint being optimized on our training set data. The hyperparemeter space was searched using a combination of random search, grid search, Bayesian optimization, and human-guided search. Once an optimal set of hyperparameters were identified, a final RF model was trained using all projections from the training set as inputs and evaluated on projections from the test set, using the AUROC as the primary measure of accuracy.
Once the primary RF model had been trained and evaluated, we trained another secondary models using the same methods as above as a comparator for our primary model. This model was trained on re-sampled instances from the evaluation set and used alternative positive labels indicating general ischemic stroke as inputs. The resulting model was then trained to predict the 10 year likelihood of general ischemic stroke.
Latent causes
The data generation process of this problem of recognizing latent causes of CS in the health record is displayed in Fig. 3. The disease process, along with its causes and effects, are the most upstream latent nodes on the causal graph. These cause a set of observations, $O$, which are recorded into the EHR by the clinician. These observations include diagnosis codes, demographics, laboratory values, and medication records, and they may influence one aother. Then, based solely on the diagnosis code observations in a given record, a binary yes/no label for CS can be applied by the process detailed in Fig. 2.
[…]
(a) Ideal causal graph where conceptual latent variables are separated into individual nodes. Node $O$ represents a network of many EHR observations which may influence each other.
[…]
(b) Causal graph where the observation network $O$ from Fig.3(a) is expanded into an arbitrary example arrangement of 4 observation nodes, each with its own error term identifiable by ICA. These error terms are the root causes of our label to the degree that the data permits for identification.
Figure 3: Two views of the structural causal graph for an arbitrary patient. The green box corresponds to the binary yes/no CS label, shaded blue nodes correspond to observations in the EHR, and dashed lines represent latent nodes.
The true root causes of our final label exist within the parentless ”Causes” node in Fig. 3(a). However, we only have access to imperfect observations of the effects of these causes [25, 27]. Under the LiNGAM model assumptions [36], the latent sources whose expressions $\mathbf{s}^{\prime}_{j}$ are captured by our ICA source matrix S represent unobserved latent causes of the target sink node (in this case, our CS label)[27, 26]. Formally, these disentangled sources correspond to the exogenous independent error terms of the structural equation model that we assume to describe the underlying causal process of our problem (Fig. 3(b)). Since causal relations are transitive, these error terms do in fact exert a causal relationship onto the final label assigned to each record. In summary, these error terms (dashed nodes in Fig. 3(b)) are compact representations of the true unobserved causes inferred using the data available in the EHR; therefore, the data generative process in Fig. 3(a) can be equivalently be represented by Fig. 3(b)[27].
Feature importance
To compare the degree of causation each source has on CS in our migraine patients, we computed the SHapley Additive exPlanations (SHAP) value[37] for each model feature across each record in the test set. Because the features used by our model that predicts the probability $P(\text{label}\ |\ \textbf{S})$ are the mutually independent error terms of our structural causal model (Fig. 3(b)), the SHAP value of each input feature is a quantitative estimate of the causal effect of that latent source on the record’s final label.[27]. If the SHAP value is positive, then that source increases the probability of the positive CS label, while a negative SHAP value signifies that the source decreases the probability of a positive CS label.
Results
Model performance
The mean AUROC of our model when evaluated via 6-fold cross validation with optimal hyperparameters was 0.771. The AUROC when trained on all training set projections and evaluated against the test set was 0.782 (Fig.4(a)). This is lower than the test accuracy of our secondary RF model trained to assess sources of general ischemic stroke, which had a test AUROC of 0.805. The accuracies of both of these models are in line with other published cardiovascular risk estimators that see regular clinical use, such as the ASCVD risk score which has been shown to have an AUROC between 0.65-0.80[38, 39].
Inferred causes of CS
(a) Above: Waterfall chart showcasing the model’s decision-making for a random patient record, illustrating its ability to provide patient-level causal explanations for a record’s final label.
Below: ROC curve for test set evaluation (AUC = 0.782).
(b) Distribution of SHAP values, color-coded by source expression value, for the 10 sources found to be most predictive of cryptogenic stroke. Sources are ranked in descending order by mean absolute SHAP value (in log-odds units). Positive SHAP values indicate an increased stroke risk, while negative SHAP values indicate a reduced stroke risk. Signatures with boxed labels are uniquely causative of cryptogenic stroke (i.e. did not also rank as a top-10 source for general IS).
The majority of these sources correspond to common drug regimens used to prevent stroke; they are unsurprisingly negative predictors in the model. However, sources with dominant features of phenytoin and allergic rhinitis—which were a negative and positive predictor, respectively—were unexpected inclusions among the list of top sources.
Figure 4:
Inferred sources were arranged in descending order by their mean absolute SHAP value, with the top 10 most predictive sources shown in Fig.4(b). The majority of these sources follow a similar pattern: the highest-featured (and usually dominant) chart variable is a medication known to reduce stroke risk, while the remaining variables comprise the constellation of clinical diagnoses and symptoms one would expect to see in a patient taking said medication (Fig. 5). Records with negative expressions for these sources largely had positive SHAP values (indicating increased predicted stroke risk), and vice versa.
Figure 5: Signature description diagrams for 4 of the most predictive sources, as shown in Fig.4(b). Bar length represents the normalized change in a clinical variable’s value for each unit of expression, shown in parenthesis. Inset log-scale histograms show source expression levels across all discovery set instances.
Positive expressions of these sources are easy to interpret—they indicate that the medication is present in the record, and at least some of the features are as well. Records with a zero or near-zero value for these sources are similarly easy—neither the medication nor the associated features are present. However, records with negative expressions for these sources are harder to interpret, despite these records yielding the highest positive SHAP values for the corresponding signature. Upon chart review, records that negatively express sources are often partial matches, where some of the variables further down the signature description diagram (Fig.3) are present, while the dominant variable(s) are not. In the case of these medication-dominated sources, this looks like a record that does not contain a mention of the indicated medication but does mention diagnoses associated with said medication.
SHAP values were also computed for the general IS model, and the top features were compared between models. There was significant overlap between the top features. However, as many strokes are incorrectly labeled as cryptogenic,[8] this is to be expected, as is it plausible that different stroke types may share common causes. Sources related to lipid-lowering agents and antihypertensives had strong estimated causal effects in both models, while signatures uniquely causative of cryptogenic stroke are noted in Fig.4(b). These include a source primarily characterized by the anti-epileptic phenytoin, which was protective against cryptogenic stroke in the model, and a source related to allergic rhinitis, which had causal effects on CS.
Additionally, if signatures are ranked by max absolute SHAP value instead of mean absolute SHAP value, a source corresponding to sickle cell disease appears. This high max SHAP value indicates that this source is highly impactful when present in the record. However, the model does not consider lack of expression of this source to be protective, meaning records without a sickle cell diagnosis have a near-zero SHAP value. Since sickle cell disease is a rare diagnosis, this source’s average mean SHAP value remains low.
Meanwhile, the source related to allergic rhinitis sees the opposite effect. Its max absolute SHAP value is significantly lower than other top predictors. However, the model is incredibly consistent in how it estimates the effect of this source—positive expression invariably increases the risk of CS in the model, and its negative expression decreases risk. So although the associated change in CS risk is small (0.001 on average), it is consistently non-zero, resulting in a large mean absolute SHAP value.
While this may indicate allergic rhinitis itself exerts a causal effect on ischemic stroke, a possibility that has been previously (though infrequently) discussed in the literature[40, 41], another plausible explanation is that this source represents the unobserved downstream effects of allergic rhinitis. Allergic rhinitis is frequently treated with over-the-counter medications[42] which are often not recorded in the EHR[43]. It is possible, then, that this source represents the cumulative effect of chronic usage of sympathomimetic decongestants such as oxymetazoline and pseudoephedrine which have been linked to CS in case reports and retrospective studies.[44, 45]
Interestingly, sources relating to previous diagnoses of thrombotic or lacunar stroke appeared to be highly protective against CS, despite the fact that occurrence of a first stroke is known to increase the risk of a subsequent stroke significantly across all types[46]. This unexpected result is likely the unintended consequence of our sampling methods. Since the sampling window for all records labeled as positive only includes dates prior to the patient’s first recorded stroke, we unintentionally blinded the model to instances of recurrent strokes. This likely caused the model to assume that such events were incredibly rare or impossible, leading to the conclusion that a prior stroke must somehow prevent future cryptogenic stroke.
Conclusions and discussion
In this study, we proposed using machine learning algorithms to implement a probabilistic independence-based analysis in order to shed light into the complex relationship between cryptogenic stroke and migraine. We used ICA to disentangle 2000 independent latent sources of disease and their corresponding signatures from a large real-world EHR dataset of neurology patients. This was done unsupervised, which removes the inherent error and bias that comes with human analysis of data, allowing for the data to speak for itself. However, during signature discovery, we were limited by the memory constraints of our machine, forcing us to settle to untangle a maximum of 2000 signatures. The dimensional reduction during projection of our 9000-dimensional patient cross-sections onto the 2000-dimensional signature space implies that information is lost during this process, though it is unclear to what extent. The natural implication is that multiple sources of disease may be combined into one signature, reducing interpretability of our results.
The equivalence between the sources disentangled by ICA and the error terms in the patient-specific causal graph allows us to utilize SHAP values as a way to quantify the the causal effect each source has on each patient’s final prediction within the model. This results in increased interpretability and usefulness beyond that seen with typical predictive models. Many clinically used risk calculators can identify patients at risk for a given disease, but cannot give reasons for why that patient might be at risk. Therefore, causal models such as ours can be particularly useful when trying to interpret the prediction of individual patient records where the data is particularly sparse, incoherent, or complex.
On larger scale, this enhanced interpretability allows for the identification of the most common root causes of CS at a population level. Overall, the sources with the largest causative effects were protective and were characterized by pharmacological treatment with common preventative medications. However, two surprising drivers of cryptogenic stroke in migraine identified by the model were a source related to phenytoin, which was protective against CS, and a source related allergic rhinitis, which was causative. These differ significantly from the other sources with large causative effects and are promising directions for future investigation.
Acknowledgements
In utilizing VUMC’s Synthetic Derivative, this project was supported by CTSA award No. UL1TR000445 from the National Center for Advancing Translational Sciences. Its contents are solely the responsibility of the authors and do not necessarily represent official views of the National Center for Advancing Translational Sciences or the National Institutes of Health.